Session: Sep 1999 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Jader Monari at Sun Sep 26 12:38:17 METDST 1999) Minor failures
Power supply failure during the observations

Noto (C.Trigilio at Mon Sep 27 11:12:38 METDST 1999) Success

Torun (K.Borkowski at Mon Sep 27 14:19:30 METDST 1999) Success

Effelsberg (A. Kraus at Tue Sep 28 10:40:54 METDST 1999) Minor failures
Some scans have been lost due to insufficient slewing time. Lots of parity errors occured.

Onsala (A. Polatidis at Wed Sep 29 12:11:14 METDST 1999) Failure
Well major Failure. We lost all observations until 0800 UT due to an error in the procedure file. Only data from 0804 UT are valid and we will only send the last tape at the correlator. We are very sorry for this.

Jodrell (Alastair Gunn at Wed Sep 29 13:31:35 METDST 1999) Minor failures
PI did not give enough slew time for J0241-08 scan at 22:58 so schedule got a bit confused. Strong RFI. No other known problems.

Calibration (Jean\-Francois Desmurs at Sun Nov 7 15:53:19 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the sep99 directory under the name: ep027a.antab. Read the comments in the file to get further information (Rem : Jb overflow on VC05 and 06.)

JIVE (Bob Campbell at Wed Mar 1 08:56:57 MET 2000) Minor failures
Mc after pass9 of tape1 (2118); No On before pass7 of tape 3 (0804); No Tr throughout -- hence never more than 5 stations; mostly 4.
Jb track12 dead --> LCP 1st subband suspect; Wb tracks 7,9,10,15 marginal --> RCP 3rd subband, LCP 1st & 3rd subband could be suspect (both arising from the interplay between subband/polarization channels & their constituent tracks via fan-in).
aips++ plots given to Co-Inv. during visit to JIVE: all weight(t), amp(lag); selected amp(t), phs(t), amp(frq).